OpenStudy (jeremiah18):

What is the speed of a car in feet/second that is traveling at 60 miles/hour?

OpenStudy (jewishpriest):

(1 mile = 5280 ft.) In the previous question, we determined that 60 mph is the same as 88 ft per second. So, 50 mph would be 50/60 of 88 ft. per sec. 50/60 of 88 can be found by multiplying 50/60 or 5/6 x 88. 5/6 x 88 = 73 1/3 or 73.33 So the car is traveling an average speed of 73 1/3 ft per sec for 15 second.

OpenStudy (mathstudent55):

\(\Large 60 \dfrac{mi}{h} \times \dfrac{1 h}{60 mn} \times \dfrac{1 mn}{60 s} \times \dfrac{5280 ft}{1 mi} = 88 \dfrac{ft}{s}\)

OpenStudy (mathstudent55):

@JewishPriest The problem asks what 60 mph is in fps. If in the previous problem you found out that 60 mph is 88 fps, then why do you need all those calculations to end up with the wrong answer?